Debra usually buys a soft drink when she goes to a movie theatre, where she has a choice of three sizes: the 8-ounce drink costs $1.50, the 12-ounce drink $2.00, and the 16-ounce drink $2.25. Describe the budget constraint that Debra faces when deciding how many ounces of the drink to purchase. (Assume that Debra can costless dispose of any of the soft drink that she does not want.


1
Expert's answer
2021-09-30T12:20:44-0400

From the given information in the question we can notice that, the price per ounce is decreasing as the size of drink is increasing.

That is,

- When she buys the 8 ounce soft drink, she pays : $1.50 / 8 oz = $0.19 per ounce.

- When she buys the 12 ounce size, she pays: $2.00 / 12 oz = $0.17 per ounce.

- And when she buys the 16 ounce size, she pays: $2.50 / 16 oz = $0.14 per ounce.

So, for example, if Debra wants 16 ounces of soft drink, she should buy the 16-ounce size and not two 8-ounce size drinks. Also, if Debra wants 14 ounces, she should buy the 16-ounce drink and dispose of the last 2 ounces.

The problem assumes she can do this without cost. Given that there are three different prices per ounce of soft drink, the budget line will have two kinks in it, as illustrated in the diagram below: